# Adaptive Intelligence: leveraging insights from adaptive behavior in animals to build flexible AI systems

### v1

@@
Our world is a series of predictions made in our mind. That is, to us, the world is a construct, built conceptually from a series of millions of predictions harbored in the neural code. When we walk on the beach, we accumulate photons with our eyes, sound waves with our ears, we feel the wind on our face, and the sand shifting under our feet, and all of this comes together to give us the perception of a warm, sunny day with ocean waves crashing on the shore. However, our brain sits in the dark confines of our skull and tens of billions of neurons communicate through electrical activity in a bath of chemicals. As incoming sensory information is processed in the brain at different times, the brain imposes a lag in order to compile it, thrusting our perception into the past—tens of precious milliseconds behind our perceived reality. Nevertheless, we need to act quickly, and our brains have evidently evolved solutions to overcome these delays.
@@
Dating at least as far back as Aristotle, there has been a philosophical and scientific question of how our brains build models of the world (11). How is it that we can combine sensory information into perception? How long does this take? Are we always tens, or even hundreds of milliseconds behind reality? How could our brains make predictions about the world in order to act faster? Why do we mostly notice when the world violates our predictions? – if the ocean were suddenly quiet, we would instantly worry we had lost our hearing and not doubt that the ocean had completely changed. While this is familiar in everyday life, we also need to game-ify this approach to study it in the laboratory.
@@
In sensorimotor control there has been a long line of work on rapid learning. So-called motor adaptation studies has been developed where visual or proprioceptive information is perturbed such that within-session, an animal must adapt. This has not only spurred the development of new neural analysis tools (27, 28, 29, 15), but a host of behavioral tasks. For example, visuomotor rotations have been used in humans and non-human primate studies to study how they can account for sensorimotor discrepancies in only a few hundred trials (30, 31, 32, 27, 33). The same principle is used for changing environmental dynamics of a manipulandum that causes deviations in a limb-movement tasks (34, 35, 7, 36). Notably, these tasks are specifically designed to measure both adaptive learning, and the formation of internal models. There is always a period of baseline control movements, perturbations, and a return to baseline-condition that allows researchers to behaviorally measure if an internal model was updated (Figure 1c).
Neurally, evidence shows that motor and sensory areas can change their tuning properties during the course of learning (37, 38, 36, 39, 40). A study by Sun et al. (38) showed that during a motor adaptation task in macaques, motor cortex (M1) effectively index memories of the hand-force required for learning to adapt. Notably, the neural subspace most predictive of hand forces changed during the period before movement (preparatory), specifically during the learning epoch. In a neural dimension orthogonal to this force-predictive subspace, they identified a uniform shift across all movement directions, including those unaffected by learning. Intriguingly this uniform shift remained after exposure to the force field, reflecting an updated internal model. Other works also show new evidence of state-changes that persist across learning (36, 8), and temporally-resolved prediction errors (41, 36). Influential work on visual cortex (V1) has shown the existence of ‘mismatch’ neurons when the expected visual feedback is disrupted (41), and similar prediction errors have been found in somatosensory (S1), motor (M1), and frontal areas of cortex (40, 36).
@@
Figure 2: Neural computations: biological teaching signals. (a) Diagram of the theoretical prediction error circuit (based on evidence from Keller et al. (41), Leinweber et al. (44), Green et al. (45)). (b) Diagram of the computation in the reward prediction error circuit of the midbrain (ventral tegmental area) dopamine and GABAerigc neurons. Inspired from Gershman et al. (46), Cohen et al. (47).
Related to cell types, the anatomy of the brain is important to consider, as this too could have direct implications for the future design of adaptive AI systems. Several areas in the brain have this incredible structure, – layers in the cerebral (neo-) and cerebellar cortex – which could be a critical part of new architectures for AI systems. There is a long-standing debate about representation emergence from data vs. architecture constraints (48, 49), but it is increasingly clear that architecture is critically important and linked to function.
@@
Form ties to function, and one emerging hot topic is how cortical circuits implement learning from prediction errors across cell types and layers. Building on the initial discovery of sensory prediction errors in cortex (41), additional works have begun to record from subtypes of neurons such as parvalbumin (PV), VIP, SST, and excitatory neurons in order to form a more complete model of how both excitatory and inhibitory neurons across layers could implement the learning rule (45, 51, 44) (Figure 2a). This work in cortex follows earlier work that discovered reward prediction errors (RPEs) in the midbrain (within the ventral tegmental area) (52, 47, 46), and they are also part of an inhibitory-gated circuit, where GABAergic neurons play a critical role in the computation of RPEs (53) (Figure 2b).
@@
Recent work aims to unify a view for how a hierarchical implementation of prediction errors (including RPEs) could aid in learning across sensory, cognitive, and motor systems (54). Tsai et al. (54) et al. found that during a cue-guided reward task, layer 2/3 somatosensory neurons showed enhanced responses to reward-predictive stimuli, and this led to reduced reward-prediction errors and increased confidence in predictions. Following rule reversal, the lateral orbitofrontal cortex, via VIP interneurons, signaled context-prediction errors, reflecting a loss of confidence. The work suggests a hierarchical interaction of prediction errors across cortical regions, with top-down signals modulating sensory cortex activity Notably, prior work finds limited roles for only “pure” reward-based learning in motor learning, but it can be combined with sensory prediction errors to shape performance (32, 7, 55). I envision that future work will more deeply test how various learning signals are used concurrently across systems. For example, tasks that have different types of prediction errors can be developed that might conflict, forcing the subject to weight different errors in order to guide future actions.
@@
In the context of learning and links to AI, another critical view on cell types and anatomy is neural development. For example, cells in the cortex migrate to their final resting location in different times. Moreover, winner-takes-all synapses form, such as at the neuro-muscular-junction of alpha motor neurons and muscles. There is also a complex and highly regulated developmental program in the spinal cord of transcription factors and hox genes – the genetics lay out the global (rostral-caudal) and even local (dorsal-ventral zones with a level) connections. Moreover, minimally in cortex, early traveling, spontaneous waves form the basis of early connections between cortical neurons.
@@
Intriguingly, new work has highlighted how the transformer architecture naturally gives rise to waves, which are a critical feature in the brain (56). Waves are not only seen in early development (as mentioned), but also in steady-state – where they are thought to keep the brain’s subregions interconnected and sharing neural dynamics that could encode global states and arousal levels.
@@
New tools – i.e., the third ingredient – are critical for the study of adaptive behavior. Computational neuroscience has made large advances in the past decade, fueled by the need to better understand ensembles of neurons (17, 57, 58, 59) and by leveraging the concurrent innovations in deep learning, which is reviewed in Hurwitz et al. (60) and Mathis et al. (15). In brief, algorithms such as variational autoencoders, transformers, contrastive learning-optimized neural networks, and diffusion models are already making big strides in our ability to combine data across subjects, measure model multi-region interactions, and ‘decode the brain’ (15).
@@
For review, I want to briefly highlight several tools that enable time-series analysis in single neurons and neural populations, as these are critical for adaptive learning studies, and I suspect will become more popular tools for studying AI systems. One of the most impactful tools in the last 20 years has been the adoption of generalized linear models (GLMs) for neuroscientific applications (61, 62). Here, a model is trained using a feature matrix (can be one or many features) with the goal of predicting the spikes of neurons. While GLMs can come in many flavors, as an example, to handle variability in behavior, models are trained and tested on different splits of the data (also for cross validation). Predictions from these test sets are combined, and a pseudo-R2 is calculated. This does require that behavior is repeated, but also allows for flexibility; i.e., there is no direct trial-averaging or behavior-triggered averaging of the neural code.
@@
For ensembles of neurons, several methods have been used, ranging from PCA for measuring neural trajectories (63, 64), contrastive learning paired with consistency metrics to measure changes in the latent neural dynamics across learning (36, 29), to nonlinear dynamical systems for task- and task-irrelevant metrics (65, 66). These are more deeply covered in reviews elsewhere (60, 15), but I will note that several approaches can be used for single-trial dynamics and have been especially impactful in motor learning and adaptation studies in non-human primates (59, 67). In particular, work by Azabou et al. (68) introduce a training framework and architecture designed to model population dynamics in large-scale neural recordings, which tokenizes individual spikes to capture fine temporal structures and constructs a latent representation of neural population activity. They trained this on nearly 100 hours of data and show excellent cross-task generalizing in primate reaching studies. One limitation of this approach is the requirement of labels at test-time, but in practice this is nearly always available.
@@
I predict we are just on the cusp of an influx of new advances in (neuro)science due to advances in AI. For example, large-language models (LLMs) are already starting to appear in behavioral neuroscience (69) for digitizing and quantifying actions in video-based data. New vision-language, hierarchical, and multi-task foundation models are surely coming. For neural analysis, already several groups are actively working on building more unified encoders that can not only be used for brain-machine interfaces, but they themselves could serve as models of the brain (68, 29, 70, 71, 15). This is sure to give rise to new approaches to extract meaningful computational principles from neural dynamics, and these new learning rules have great potential to directly influence how we train, optimize, and deploy machine learning systems.
@@
Continual learning aims to address the challenge of developing models that can learn incrementally from a continuous stream of data without forgetting previously acquired knowledge. Traditional machine learning models often face catastrophic forgetting (72) when retrained on new data, but new advancements have introduced new techniques like elastic weight consolidation (73, 74), synaptic intelligence, and memory-replay (75, 76) to mitigate these issues.
While not directly brain-inspired, Elastic Weight Consolidation (EWC) works by constraining certain parameters to reduce interference between tasks, effectively mitigating catastrophic forgetting (73). More recent work as been adapting EWC to fine-tuned self-supervised models, which has yielded performance improvements on tasks that have biased datasets by maintaining knowledge of previous tasks better than traditional methods (74, 77). If we attempt to map this to neural dynamics, this might be related to the specialization of circuits. Namely, the “parameters” of certain areas could be fixed (such as primary receptors), while others could be more plastic (such as hippocampus and neocortex).
@@
Another brain-inspired machine learning algorithm has been the development of memory-replay (79, 80), which mimics the memory systems in the hippocampus (Figure 3a).  [^2] Notably, memory-replay also, perhaps indirectly, build on the memory-inspired system of the Hopfield Network, which is a type recurrent neural network designed for content-addressable memory, resembling a spin glass system (81), whose invention also won John Hopfield a Nobel Prize in 2024. Memory-replay aims to have a memory bank of already-learned actions (or images or tokens, etc.) that can be “replayed” (interweaved) into the training data in order to limit catastrophic forgetting (Figure 3b). This became a popular technique in reinforcement learning, and more recently, in computer vision and in LLMs. This training approach can also be used during active learning when new data (unlabeled, pseudo-labeled or labeled) can be used to fine-tune a model. If the input data continues to morph out-of-distribution (OOD) then using memory-replay can be leveraged to be sure the prior performance can be achieved.
@@
Lastly, one of the most biologically inspired and grounded advances has been in the development of spiking neural networks (SNNs) in the 1990’s. Called the “third generation of neural networks” (86), SNNs are unique in their ability to model a time-dependent spike process, making them not only biologically more plausible, but also highly energy efficient and can be used on edge-computing devices. Notably, SNNs naturally lend themselves to hardware accelerated devices and neuromorphic computing. Platforms such as Intel’s Loihi, IBM’s TrueNorth, and SpiNNaker provide hardware specifically designed to run SNNs efficiently (87). However, historically they have been difficult to train and scale, but that’s rapidly changing and new methods to transfer information from ANNs to SNNs are emerging  (88).
@@
While many questions remain about how these models are implemented in neural dynamics, there is increasing evidence of core computations (such as prediction errors) across brain regions, and there are architectural constraints on these systems that stem from brain-body dynamics. Brain regions are often based on primary sensors: vision, olfaction, touch, proprioception, audition, and so forth. Yet, as Mountcastle argued, it’s less about a “sensory” or “motor” space, but rather a function of inputs and outputs, and by extension, inter-connectivity of regions. Thus, as I will argue below, we should take inspiration from building specialized nodes that have core computations in the right reference framework that are interconnected in order to act.
@@
Thus, I argue we should also focus on building agents that are imbued with adaptive algorithms found in animal intelligence, and not only build unified monolithic Foundation Models. There is emerging work on artificial agents that I believe will make this possible. Agents that act as operation systems are one promising avenue. They can be a series of LLMs (which is currently being explored (82, 69)), but also a series of VLMs  (92), or jointly optimized encoders that individually are first pretrained (think mini-Foundation Models, such as an encoder for animal behavior (76) or unified neural region-specific pretrained encoders (68)). Then, these series of models could leverage the ‘learning’ techniques discussed above, such as memory-replay or synaptic intelligence. This is also highly likely to drive innovations, as it is also not clear that having many, independent losses is optimal.
@@
How could this work? Imagine we have a series of core-expert models (aka nodes): one for segmenting video, one for extracting poses, another for ingesting text plus images (VLMs), and one for extracting dynamics from neural spikes (Figure 4c). Each could be first experts in their specific domain, then linked via a shared latent space (even in a common spike-like framework) and optimized together. This could have a human-level language interaction agent (such as a LLM) where at inference-time the multi-modal input data could also be queried with natural language, but importantly, monitor the sub-modules (much like a prefrontal cortex). Then, we could dynamically “lock” models that show robustness on out-of-distribution samples and robustness to adversarial attacks. These nodes don’t then need to be trained (adapted) until they are deemed no longer robust. This proposal would be akin to the cortex–basal ganglia loop that has shown to be instrumental in learning and habit formation (93). Namely, these nodes could be in states of “skill learning” or “habituation” (frozen) and be constantly monitored for robustness (Figure 4c). When robust levels fall, then continual learning, memory-replay, pseudo-labeling or injection of new high-quality labeled data could be added without taking the entire system offline. There has been a surge of exciting work on building OOD detection modules (94, 95) and explainable attribution methods in time-series and images (96, 97, 98, 99, 100) that could make this an attractive path forward.
@@
Another approach, closely linked to robustness detection, could be specifically adding in prediction errors into the LLMs. Effectively, prediction errors give a signed teaching signal when an OOD sample is detected. Early works on predictive-coding inspired networks already showed promise (101), and now with modern transformer architectures or hybrid CNN-transformers, perhaps we can leverage prediction errors in order to do not only in-context learning, but signal which nodes need a new model – akin to how the brain must decide to online change a motor command vs. update an internal model. Here, we can take inspiration from the motor system (102). A one-time error means “in-context learn” and change your ongoing motor command [^3] While neurally it remains to be seen if such rapid learning (on the order of ms) can cause synaptic changes, it is doubtful that even behaviorally-time-dependent-plasticity would be fast enough., but a repetitive error means update your internal model (Figure 1c). If the incoming perturbation cannot be solved with online learning or model adaptation, a new internal model would then need to be created (which takes more time and can be completed “offline”). The neural schema in prefrontal cortex discovered by El-Gaby et al. (20) could inspire new approaches for adaptive memory-replay and new model building as well (Figure 1b).

### v2

@@
Our world is built on a series of predictions made in our mind. That is, to us, the world is a construct, created conceptually from a series of millions of predictions harbored in the neural code. When we walk on the beach, we accumulate photons with our eyes, sound waves with our ears, we feel the wind on our face, and the sand shifting under our feet, and all of this comes together to give us the perception of a warm, sunny day with ocean waves crashing on the shore. However, our brain sits in the dark confines of our skull and tens of billions of neurons communicate through electrical activity in a bath of chemicals. As incoming sensory information is processed in the brain at different times, the brain imposes a lag in order to compile it, thrusting our perception into the past—tens of precious milliseconds behind reality. Nevertheless, we need to act quickly, and our brains have evidently evolved solutions to overcome these delays.
@@
Dating at least as far back as Aristotle, there has been a philosophical and scientific question of how our brains build models of the world (11). How is it that we can combine sensory information into perception? How long does this take? Are we always tens, or even hundreds of milliseconds behind reality? How could our brains make predictions about the world in order to act faster? Why do we mostly notice when the world violates our predictions? – if the ocean were suddenly quiet, we would instantly worry we had lost our hearing and not doubt that the ocean had completely changed. While this is familiar in everyday life, we need to game-ify this approach to study it in the laboratory.
@@
In sensorimotor control there has been a long line of work on rapid learning. So-called motor adaptation studies have been developed where visual or proprioceptive information is perturbed such that within-session, an animal must adapt. This has not only spurred the development of new neural analysis tools (27, 28, 29, 15), but a host of behavioral tasks. For example, visuomotor rotations have been used in humans and non-human primate studies to study how they can account for sensorimotor discrepancies in only a few hundred trials (30, 31, 32, 27, 33). The same principle is used for changing environmental dynamics of a manipulandum that causes deviations in a limb-movement tasks (34, 35, 7, 36). Notably, these tasks are specifically designed to measure both adaptive learning, and the formation of internal models. There is always a period of baseline control movements, perturbations, and a return to baseline-condition that allows researchers to behaviorally measure if an internal model was updated (Figure 1c).
Neurally, evidence shows that motor and sensory areas can change their tuning properties during the course of learning (37, 38, 36, 39, 40). A study by Sun et al. (38) showed that during a motor adaptation task in macaques, motor cortex (M1) can effectively index memories of the hand-force required for learning to adapt. Notably, the neural subspace most predictive of hand forces changed during the period before movement (preparatory), specifically during the learning epoch. In a neural dimension orthogonal to this force-predictive subspace, they identified a uniform shift across all movement directions, including those unaffected by learning. Intriguingly this uniform shift remained after exposure to the force field, reflecting an updated internal model. Other works also show new evidence of state-changes that persist across learning (36, 8), and temporally-resolved prediction errors (41, 36). Influential work on visual cortex (V1) has shown the existence of ‘mismatch’ neurons when the expected visual feedback is disrupted (41), and similar prediction errors have been found in somatosensory (S1), motor (M1), and frontal areas of cortex (40, 36).
@@
Figure 2: Neural computations: biological teaching signals. (a) Diagram of the theoretical prediction error circuit (based on evidence from Keller et al. (41), Leinweber et al. (44), Green et al. (45), Jordan and Keller (46)). (b) Diagram of the computation in the reward prediction error circuit of the midbrain (ventral tegmental area) dopamine and GABAerigc neurons. Inspired from Gershman et al. (47), Cohen et al. (48).
Related to cell types, the anatomy of the brain is important to consider, as this too could have direct implications for the future design of adaptive AI systems. Several areas in the brain have this incredible structure, – layers in the cerebral (neo-) and cerebellar cortex – which could be a critical part of new architectures for AI systems. There is a long-standing debate about representation emergence from data vs. architecture constraints (49, 50), but it is increasingly clear that architecture is critically important and linked to function.
@@
Form ties to function, and one emerging hot topic is how cortical circuits implement learning from prediction errors across cell types and layers. Building on the initial discovery of sensory prediction errors in cortex (41), additional works have begun to record from subtypes of neurons such as parvalbumin (PV), VIP, SST, and excitatory neurons in order to form a more complete model of how both excitatory and inhibitory neurons across layers could implement the learning rule (45, 52, 44) (Figure 2a). This work in cortex follows earlier work that discovered reward prediction errors (RPEs) in the midbrain (within the ventral tegmental area) (53, 48, 47), and they are also part of an inhibitory-gated circuit, where GABAergic neurons play a critical role in the computation of RPEs (54) (Figure 2b).
@@
Recent work aims to unify a view for how a hierarchical implementation of prediction errors (including RPEs) could aid in learning across sensory, cognitive, and motor systems (55). Tsai et al. (55) et al. found that during a cue-guided reward task, layer 2/3 somatosensory neurons showed enhanced responses to reward-predictive stimuli, and this led to reduced reward-prediction errors and increased confidence in predictions. Following rule reversal, the lateral orbitofrontal cortex, via VIP interneurons, signaled context-prediction errors, reflecting a loss of confidence. The work suggests a hierarchical interaction of prediction errors across cortical regions, with top-down signals modulating sensory cortex activity Notably, prior work finds limited roles for only “pure” reward-based learning in motor learning, but it can be combined with sensory prediction errors to shape performance (32, 7, 56). I envision that future work will more deeply test how various learning signals are used concurrently across systems. For example, tasks that have different types of prediction errors can be developed that might conflict, forcing the subject to weight different errors in order to guide future actions.
@@
In the context of learning and links to AI, another critical view on cell types and anatomy is neural development. For example, cells in the cortex migrate to their final resting location in different times. Moreover, winner-takes-all synapses form, such as at the neuro-muscular-junction of alpha motor neurons and muscles. There is also a complex and highly regulated developmental program in the spinal cord of transcription factors and hox genes – the genetics lay out the global (rostral-caudal) and even local (dorsal-ventral zones with a level) connections. Additionally, there is the so-called genomic bottleneck to consider, which aids in determining early innate abilities of animals – from the immediate suckling reflex to the slightly more time-intensive but still impressive ability of a horse to stand and walk a few hours after birth (57).
@@
Moreover, minimally in cortex, early traveling spontaneous waves form the basis of early connections between cortical neurons. Intriguingly, new work has highlighted how the transformer architecture naturally gives rise to waves, which are a critical feature in the brain (58). Waves are not only seen in early development (as mentioned), but also in steady-state – where they are thought to keep the brain’s subregions interconnected and sharing neural dynamics that could encode global states and arousal levels.
@@
New tools – i.e., the third ingredient – are critical for the study of adaptive behavior. Computational neuroscience has made large advances in the past decade, fueled by the need to better understand ensembles of neurons (17, 59, 60, 61) and by leveraging the concurrent innovations in deep learning, which is reviewed in Hurwitz et al. (62) and Mathis et al. (15). In brief, algorithms such as variational autoencoders, transformers, contrastive learning-optimized neural networks, and diffusion models are already making big strides in our ability to combine data across subjects, measure model multi-region interactions, and ‘decode the brain’ (15).
@@
For review, I want to briefly highlight several tools that enable time-series analysis in single neurons and neural populations, as these are critical for adaptive learning studies, and I suspect will become more popular tools for studying AI systems. One of the most impactful tools in the last 20 years has been the adoption of generalized linear models (GLMs) for neuroscientific applications (63, 64). Here, a model is trained using a feature matrix (can be one or many features) with the goal of predicting the spikes of neurons. While GLMs can come in many flavors, as an example, to handle variability in behavior, models are trained and tested on different splits of the data (also for cross validation). Predictions from these test sets are combined, and a pseudo-R2 is calculated. This does require that behavior is repeated, but also allows for flexibility; i.e., there is no direct trial-averaging or behavior-triggered averaging of the neural code.
@@
For ensembles of neurons, several methods have been used, ranging from PCA for measuring neural trajectories (65, 66), contrastive learning paired with consistency metrics to measure changes in the latent neural dynamics across learning (36, 29), to nonlinear dynamical systems for task- and task-irrelevant metrics (67, 68). These are more deeply covered in reviews elsewhere (62, 15), but I will note that several approaches can be used for single-trial dynamics and have been especially impactful in motor learning and adaptation studies in non-human primates (61, 69). In particular, work by Azabou et al. (70) introduce a training framework and architecture designed to model population dynamics in large-scale neural recordings, which tokenizes individual spikes to capture fine temporal structures and constructs a latent representation of neural population activity. They trained this on nearly 100 hours of data and show excellent cross-task generalizing in primate reaching studies. One limitation of this approach is the requirement of labels at test-time, but in practice this is nearly always available.
@@
I predict we are just on the cusp of an influx of new advances in (neuro)science due to advances in AI. For example, large-language models (LLMs) are already starting to appear in behavioral neuroscience (71) for digitizing and quantifying actions in video-based data. New vision-language, hierarchical, and multi-task foundation models are surely coming. For neural analysis, already several groups are actively working on building more unified encoders that can not only be used for brain-machine interfaces, but they themselves could serve as models of the brain (70, 29, 72, 73, 15). This is sure to give rise to new approaches to extract meaningful computational principles from neural dynamics, and these new learning rules have great potential to directly influence how we train, optimize, and deploy machine learning systems.
@@
Continual learning aims to address the challenge of developing models that can learn incrementally from a continuous stream of data without forgetting previously acquired knowledge. Traditional machine learning models often face catastrophic forgetting (74) when retrained on new data, but new advancements have introduced new techniques like elastic weight consolidation (75, 76), synaptic intelligence, and memory-replay (77, 78) to mitigate these issues.
While not directly brain-inspired, Elastic Weight Consolidation (EWC) works by constraining certain parameters to reduce interference between tasks, effectively mitigating catastrophic forgetting (75). More recent work as been adapting EWC to fine-tuned self-supervised models, which has yielded performance improvements on tasks that have biased datasets by maintaining knowledge of previous tasks better than traditional methods (76, 79). If we attempt to map this to neural dynamics, this might be related to the specialization of circuits. Namely, the “parameters” of certain areas could be fixed (such as primary receptors), while others could be more plastic (such as hippocampus and neocortex).
@@
Another brain-inspired machine learning algorithm has been the development of memory-replay (81, 82), which mimics the memory systems in the hippocampus (Figure 3a).  [^2] Notably, memory-replay also, perhaps indirectly, build on the memory-inspired system of the Hopfield Network, which is a type recurrent neural network designed for content-addressable memory, resembling a spin glass system (83), whose invention also won John Hopfield a Nobel Prize in 2024. Memory-replay aims to have a memory bank of already-learned actions (or images or tokens, etc.) that can be “replayed” (interweaved) into the training data in order to limit catastrophic forgetting (Figure 3b). This became a popular technique in reinforcement learning, and more recently, in computer vision and in LLMs. This training approach can also be used during active learning when new data (unlabeled, pseudo-labeled or labeled) can be used to fine-tune a model. If the input data continues to morph out-of-distribution (OOD) then using memory-replay can be leveraged to be sure the prior performance can be achieved.
@@
Lastly, one of the most biologically inspired and grounded advances has been in the development of spiking neural networks (SNNs) in the 1990’s. Called the “third generation of neural networks” (88), SNNs are unique in their ability to model a time-dependent spike process, making them not only biologically more plausible, but also highly energy efficient and can be used on edge-computing devices. Notably, SNNs naturally lend themselves to hardware accelerated devices and neuromorphic computing. Platforms such as Intel’s Loihi, IBM’s TrueNorth, and SpiNNaker provide hardware specifically designed to run SNNs efficiently (89). Moreover, several early work highlights their modularity and utility for multi-area modeling (90), and how multi-area SNNs could be used with semantic pointers for multi-area weighting of information (91). However, historically they have been difficult to train and scale, but that’s rapidly changing and new methods to transfer information from ANNs to SNNs are emerging  (92).
@@
While many questions remain about how these models are implemented in neural dynamics, there is increasing evidence of core computations (such as prediction errors (94, 41)) across brain regions, and there are architectural constraints on these systems that stem from brain-body dynamics. Brain regions are often based on primary sensors: vision, olfaction, touch, proprioception, audition, and so forth. Yet, as Mountcastle argued, it’s less about a “sensory” or “motor” space, but rather a function of inputs and outputs, and by extension, inter-connectivity of regions. Thus, as I will argue below, we should take inspiration from building specialized nodes that have core computations in the right reference framework that are interconnected in order to act.
@@
Thus, I argue we should also focus on building agents that are imbued with adaptive algorithms found in animal intelligence, and not only build unified monolithic Foundation Models. There is emerging work on artificial agents that I believe will make this possible. Agents that act as operation systems are one promising avenue. They can be a series of LLMs (which is currently being explored (84, 71)), but also a series of VLMs  (97), or jointly optimized encoders that individually are first pretrained (think mini-Foundation Models, such as an encoder for animal behavior (78) or unified neural region-specific pretrained encoders (70)). Then, these series of models could leverage the ‘learning’ techniques discussed above, such as memory-replay or synaptic intelligence. This is also highly likely to drive innovations, as it is also not clear that having many, independent losses is optimal.
@@
How could this work? Imagine we have a series of core-expert models (aka nodes): one for segmenting video, one for extracting poses, another for ingesting text plus images (VLMs), and one for extracting dynamics from neural spikes (Figure 4c). Each could be first experts in their specific domain, then linked via a shared latent space (even in a common spike-like framework) and optimized together. This could have a human-level language interaction agent (such as a LLM) where at inference-time the multi-modal input data could also be queried with natural language, but importantly, monitor the sub-modules (much like a prefrontal cortex). Then, we could dynamically “lock” models that show robustness on out-of-distribution samples and robustness to adversarial attacks. These nodes don’t then need to be trained (adapted) until they are deemed no longer robust. This proposal would be akin to the cortex–basal ganglia loop that has shown to be instrumental in learning and habit formation (98). Namely, these nodes could be in states of “skill learning” or “habituation” (frozen) and be constantly monitored for robustness (Figure 4c). When robust levels fall, then continual learning, memory-replay, pseudo-labeling or injection of new high-quality labeled data could be added without taking the entire system offline. There has been a surge of exciting work on building OOD detection modules (99, 100) and explainable attribution methods in time-series and images (101, 102, 103, 104, 105) that could make this an attractive path forward.
@@
Another approach, closely linked to robustness detection, could be specifically adding in prediction errors into the LLMs. Effectively, prediction errors give a signed teaching signal when an OOD sample is detected. The development of predictive-coding inspired networks already show promise (106, 107, 108), and now with modern transformer architectures or hybrid CNN-transformers, perhaps we can leverage prediction errors in order to do not only in-context learning, but signal which nodes need a new model – akin to how the brain must decide to online change a motor command vs. update an internal model. Here, we can take inspiration from the motor system (109). A one-time error means “in-context learn” and change your ongoing motor command [^3] While neurally it remains to be seen if such rapid learning (on the order of ms) can cause synaptic changes, it is doubtful that even behaviorally-time-dependent-plasticity would be fast enough., but a repetitive error means update your internal model (Figure 1c). If the incoming perturbation cannot be solved with online learning or model adaptation, a new internal model would then need to be created (which takes more time and can be completed “offline”). The neural schema in prefrontal cortex discovered by El-Gaby et al. (20) could inspire new approaches for adaptive memory-replay and new model building as well (Figure 1b).